# Corresponding Author

### v3

@@ 8.4 Multimodal Autoregressive Frameworks
Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation. Models like Emu3 wang2024emu3 demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training. This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps. Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations observed in diffusion-based PCS.
@@ 8.5 Interactive Personalization Workflow
The evolution of interactive generation systems presents valuable opportunities for PCS, particularly in addressing core challenges like the overfitting problem while enhancing user control. Recent advancements in conversational PCS systems, such as Gemini-2.0-flash’s interactive image editing team2023gemini , demonstrate how iterative refinement through natural language can optimize both subject fidelity and prompt alignment. This aligns with PCS’s fundamental goal of balancing precise subject representation with flexible editing capability.
@@ 9 Conclusion
Textual Inversion 1gal2022image Object TTF LDM Token Embedding Fine-tuning
@@
DreamBooth 2ruiz2023dreambooth Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
DreamArtist 68dong2022dreamartist Object TTF LDM Negative Prompt Fine-tuning
DVAR 69voronov2024loss Object TTF SD 1.5 Randomness Erasing
HiPer 33han2023highly Object TTF SD Token Embedding Enhancement
P+ 23voynov2023p+ Object TTF SD 1.4 Token Embedding Enhancement
Unet-finetune 72xiang2023closer Object TTF SD Parameter-efficient Fine-tuning
Jia et al. 24jia2023taming Object TTF Imagen Mask-assisted Generation; Regularization
COTI 3yang2023controllable Object TTF SD 2.0 Data Augmentation
Gradient-Free TI 73fei2023gradient Object TTF SD Evolutionary Strategy
PerSAM 21zhang2023personalize Object TTF SD Mask-assisted Generation
DisenBooth 18chen2023disenbooth Object TTF SD 2.1 Regularization
NeTI 61alaluf2023neural Object TTF SD 1.4 Token Embedding Enhancement
Prospect 30zhang2023prospect Object TTF SD 1.4 Token Embedding Enhancement
@@
Break-a-Scene 74avrahami2023break Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
COMCAT 75xiao2023comcat Object TTF SD 1.4 Attention-based Operation
ViCo 20hao2023vico Object TTF SD Attention-based Operation; Regularization
OFT 59qiu2024controlling Object TTF SD 1.5 Finetuning Strategy
LyCORIS 76yeh2023navigating Object TTF SD 1.5 Attention-based Operation
He et al. 57he2023data Object TTF SD, SDXL Data Augmentation
DIFFNAT 78roy2023diffnat Object TTF SD Regularization
MATTE 79agarwal2023image Object; Style; High-level Semantic TTF SD Regularization
LEGO 45motamed2023lego Object TTF SD Regularization
Catversion 80zhao2023catversion Object TTF SD 1.5 Embedding Concatenation
@@
CLiC 81safaee2023clic Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
@@
HiFi Tuner 48wang2023hifi Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
InstructBooth 56chae2023instructbooth Object TTF SD 1.5 Reinforcement Learning
DETEX 85cai2023decoupled Object TTF SD 1.4 Mask-assisted Generation
DreamDistribution 114zhao2023dreamdistribution Object TTF SD 2.1 Token Embedding Enhancement
DreamTuner 86hua2023dreamtuner Object TTF SD Attention-based Operation
Xie et al. 51lu2024object Object TTF SD Regularization; Mask-assisted Generation
Infusion 157zeng2024infusion Object TTF SD 1.5 Attention-based Operation
Pair Customization 160jones2024customizing Object TTF SDXL Disentanglement Approach
DisenDreamer 171chen2024disendreamer Object TTF SD 2.1 Disentanglement Approach
DreamBooth+++++ + 219fan2024dreambooth Object TTF SD Attention-based Operation
CoRe 218wu2024core Object TTF SD 2.1 Regularization
DCI 216jin2025customized Object TTF SD Token Embedding Optimization
Re-Imagen 67chen2022re Object PTA Imagen Retrieval-augmented Paradigm
Versatile Diffusion 109xu2023versatile Object PTA SD 1.4 Architecture Design
Tuning-Encoder 12gal2023encoder Object PTA SD Regularization
@@
ELITE 19wei2023elite Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
UMM-Diffusion 71ma2023unified Object PTA SD 1.5 Reference Feature Injection
SuTI 11chen2024subject Object PTA Imagen Data Augmentation
InstantBooth 22shi2023instantbooth Object PTA SD 1.4 Patch Feature Extraction
@@
BLIP-Diffusion 10li2024blip Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
Domain-Agnostic 17arar2023domain Object PTA SD Regularization; Attention-based Operation
IP-Adapter 40ye2023ip Object PTA SD 1.5 Reference Feature Injection
Kosmos-G 77pan2023kosmos Object; Face PTA SD 1.5 Multimodal Language Modeling
Kim et al. 47kim2023personalized Object PTA SD 1.4 Reference Feature Injection
CAFÉ 84zhou2023customization Object; Face PTA SD 2.0 Multimodal Language Modeling
SAG 99pan2023towards Object; Face; Style PTA SD 1.5 Gradient-based Guidance
BootPIG 41purushwalkam2024bootpig Object PTA SD Data Augmentation
CustomContrast 217chen2024customcontrast Object PTA SD 1.5 Contrastive Learning
MoMA song2024moma Object PTA SD 1.5 Multimodal LLM Adapter
@@
Diptych Prompting 209shin2024large Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
@@
StyleDrop 5sohn2024styledrop Style; Multiple Subjects TTF Muse Data Augmentation
StyleBoost 53park2023styleboost Style TTF SD 1.5 Regularization Dataset
StyleAligned 98hertz2023style Style TTF SDXL Regularization
GAL 151zhang2024generative Style; Object TTF SD 1.5 Data Augmentation
StyleForge 153park2024text Style TTF SD 1.5 Parameter-efficient Fine-tuning
FineStyle 213zhangfinestyle Style TTF Muse Mask-assisted Generation
Style-Friendly 211choi2024style Style TTF FLUX-dev; SD 3.5 SNR Sampler
StyleAdapter 26wang2023styleadapter Style PTA SD Data Augmentation; Reference Feature Fxtraction
ArtAdapter 42chen2023artadapter Style PTA SD 1.5 Data Augmentation
ProFusion 55zhou2023enhancing Face TTF SD 2.0 Fusion Sampling
Celeb Basis 60yuan2023inserting Face TTF SD Face Feature Representation
HyperDreamBooth 58ruiz2023hyperdreambooth Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
Banerjee et al. 39banerjee2023identity Face TTF SD 1.4 Regularization
Magicapture 90hyung2023magicapture Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
Concept-centric 35cao2023concept Face TTF SD 1.5 Modified Classifier-free Guidance
Cross Initialization 66pang2023cross Face TTF SD 2.1 Regularization
@@
OMG 115kong2024omg Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
@@
InstantFamily 161kim2024instantfamily Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
PersonaMagic 203li2024personamagic Face TTF SD 1.4 Attention-based Operation
Su et al. 49su2023identity Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
@@
FastComposer 54xiao2023fastcomposer Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
Face0 88valevski2023face0 Face PTA SD 1.4 Fusion Sampling
DreamIdentity 89chen2023dreamidentity Face PTA SD 2.1 Attention-based Operation
@@
Face-Diffuser 52wang2023high Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
W-Plus-Adapter 91li2023stylegan Face PTA SD 1.5; StyleGAN Face Feature Representation
Portrait Diffusion 32liu2023portrait Face PTA SD 1.5 Mask-assisted Generation
RetriNet 92tang2023retrieving Face PTA SD Mask-assisted Generation
FaceStudio 31yan2023facestudio Face PTA SD Attention-based Operation
PVA 27xu2024personalized Face PTA LDM Mask-assisted Generation
DemoCaricature 94chen2023democaricature Face; Extra Conditions PTA SD 1.5 Regularization
@@
PhotoMaker 28li2023photomaker Face PTA SDXL Data Collection; Reference Feature Extraction
Stellar 95achlioptas2023stellar Face PTA SDXL Evaluation Prompts and Metrics
@@
PortraitBooth 96peng2023portraitbooth Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
InstantID 37wang2024instantid Face; Extra Conditions PTA SDXL Multi-feature Injection
ID-Aligner 162chen2024id Face PTA SD 1.5; SDXL Feedback Learning
@@
MoA 164ostashev2024moa Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
IDAdapter 166cui2024idadapter Face PTA SD 2.1 Mixed Facial Features
Infinite-ID 169wu2024infinite Face PTA SDXL Reference Feature Injection
Face2Diffusion 170shiohara2024face2diffusion Face PTA SD Multi-feature Injection
FreeCure 210cai2024foundation Face PTA SD 1.5; SDXL Mask-assisted Generation
Omni-ID 205qian2024omni Face PTA FLUX Face representation enhancement
@@
Custom Diffusion 6kumari2023multi Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
Cones 4liu2023cones Multiple Subjects TTF SD 1.4 Concept Neurons Activation
SVDiff 8han2023svdiff Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
Perfusion 7tewel2023key Multiple Subjects TTF SD 1.5 Regularization
Mix-of-Show 46gu2024mix Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
Cones-2 62liu2023cones Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
PACGen 34li2023generate Multiple Subjects TTF SD 1.4 Mask-assisted Generation
Compositional Inversion 9zhang2023compositional Multiple Subjects TTF SD Attention-based Operation; Regularization
EM-style 44rahman2024visual Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
MC2 154jiang2024mc Multiple Subjects TTF SD 1.5 Attention-based Operation
MultiBooth 156zhu2024multibooth Multiple Subjects TTF SD 1.5 Mask-assisted Generation
Matsuda et al. 165matsuda2024multi Multiple Subjects TTF LDM Mask-assisted Generation
MagicTailor zhou2024magictailor Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
AnyDoor 14chen2023anydoor Multiple Subjects PTA SD 2.1 Mask-assisted Generation
Subject-Diffusion 16ma2023subject Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
CustomNet 29yuan2023customnet Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
MIGC 174zhou2024migc Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
Emu2 223sun2024generative Multiple Subjects PTA SDXL In-Context Learning
SSR-Encoder 221zhang2024ssr Multiple Subjects PTA SD 1.5 Attention-based Operation
$\lambda$-eclipse 222patel2024lambda Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
MS-Diffusion 220personalizationms Multiple Subjects PTA SD Attention-based Operation
ReVersion 63huang2023reversion High-level Semantic TTF SD Regularization
Inv-ReVersion 159zhang2024inv High-level Semantic TTF SD 1.5 Regularization
CusConcept 215xu2024cusconcept High-level Semantic TTF SD 2.1 Regularization
ADI 43huang2023learning High-level Semantic TTF SD 2.1 Mask-assisted Generation
PhotoSwap 15gu2024photoswap Extra Conditions TTF SD 2.1 Attention-based Operation
PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
Layout-Control 152chen2024training Extra Conditions TTF SD Attention-based Operation
SwapAnything 155gu2024swapanything Extra Conditions TTF SD 2.1 Mask-assisted Generation
PE-VITON 38zhang2023two Extra Conditions TTF SD Shape and Texture Control
Viewpoint Control 163kumari2024customizing Extra Conditions TTF SDXL 3D Feature Incorporation
Prompt-Free Diffusion 110xu2023prompt Extra Conditions PTA SD 2.0 Reference Feature Injection
Uni-ControlNet 111zhao2024uni Extra Conditions PTA SD Multi-feature Injection
ViscoNet 93cheong2023visconet Extra Conditions PTA SD Multi-feature Injection
Context Diffusion 112najdenkoska2023context Extra Conditions PTA LDM Multi-feature Injection
FreeControl 113mo2023freecontrol Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
Li et al. 168li2024tuning Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
Table 5: Paper summary on personalization.
Paper Scope Framework Technique Backbone
Tune-A-Video 150wu2023tune Video TTF SD 1.4 Parameter-efficient Fine-tuning
Gen-1 148esser2023structure Video TTF LDM Multi-feature Injection
Make-A-Protagonist 135zhao2023make Video TTF LDM Attention-based Operation
Animate-A-Story 136he2023animate Video TTF SD Retrieval-augmented Paradigm
MotionDirector 101zhao2023motiondirector Video TTF ZeroScope; ModelScope Disentanglement Approach
LAMP 102wu2023lamp Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
VMC 140jeong2023vmc Video TTF Show-1 Parameter-efficient Fine-tuning
SAVE 103song2023save Video TTF VDM Attention-based Operation
Customizing Motion 104materzynska2023customizing Video TTF ZeroScope Regularization
DreamVideo 142wei2023dreamvideo Video TTF ModelScope Disentanglement Approach
MotionCrafter 106zhang2023motioncrafter Video TTF VDM Disentanglement Approach
Customize-A-Video 172ren2024customize Video TTF ModelScope Multi-stage Refinement
Magic-Me 173ma2024magic Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
CustomTTT 202bi2024customttt Video TTF VDM Fine-tuning and Alignment
PersonalVideo 208li2024personalvideo Video; Face TTF AnimateDiff Identity Injection
@@
CustomVideo 145wang2024customvideo Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
VideoDreamer 138chen2023videodreamer Video; Multiple Subjects TTF SD 2.1 Data Augmentation
VideoAssembler 82zhao2023videoassembler Video PTA VidRD Reference Feature Injection
VideoBooth 83jiang2023videobooth Video PTA VDM Reference Feature Injection
VideoMaker 201wu2024videomaker Video PTA VDM Reference Feature Injection
SUGAR 204zhou2024sugar Video PTA VDM Attention-based Operation
StyleCrafter 212liu2024stylecrafter Video; Style PTA SD 1.5 Reference Feature Injection
ID-Animator 158he2024id Video; Face PTA AnimateDiff Reference Feature Injection
ConsisID 207yuan2024identity Video; Face PTA CogVideoX frequency decomposition
@@
MotionCharacter 206fang2024motioncharacter Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
@@
DreaMoving 105feng2023dreamoving Video; Face; Extra Conditions PTA SD Multi-feature Injection
Magic3D 120lin2023magic3d 3D TTF Imagen; LDM Score Distillation Sampling
DreamBooth3D 121raj2023dreambooth3d 3D TTF Imagen Multi-stage Refinement
PAS 122azadi2023text 3D PTA SD Text-to-3D-Pose
StyleAvatar3D 123zhang2023styleavatar3d 3D PTA LDM Multi-view Alignment
AvatarBooth 124zeng2023avatarbooth 3D TTF SD Dual Model Fine-tuning
MVDream 126shi2023mvdream 3D TTF SD 2.1 Score Distillation Sampling
Consist3D 127ouyang2023chasing 3D TTF SD Token Embedding Enhancement
Animate124 128zhao2023animate124 3D TTF SD 1.5 Multi-stage Refinement
Dream-in-4D 130zheng2023unified 3D TTF SD 2.1 Multi-stage Refinement
TextureDreamer 133yeh2024texturedreamer 3D TTF LDM Texture Extraction
TIP-Editor 134zhuang2024tip 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
Anti-DreamBooth 13van2023anti Attack and Defense TTF SD 2.1 Perturbation Learning
Concept Censorship 50zhang2023backdooring Attack and Defense TTF SD 1.4 Trigger Injection
Huang et al. 167huang2024personalization Attack and Defense TTF SD backdoor attack
Continual Diffusion 64smith2023continual Others TTF SD Continual Learning
SVGCustomization 36zhang2023text Others TTF SD 1.5 Fine-tuning and Alignment
StitchDiffusion 25wang2024customizing Others TTF LDM Parameter-efficient Fine-tuning
MC-TI 214wang2024multi Others TTF SD 1.5 Regularization

### v4

@@ 8.4 Multimodal Autoregressive Frameworks
Recent advancements in multimodal autoregressive models present novel solutions for PCS by unifying cross-modal understanding and generation. Models like Emu3 [164] demonstrate that autoregressive architectures can natively handle image-text-video sequences through discrete tokenization and joint transformer training. This paradigm enables seamless integration of user-provided multimodal references (e.g., text descriptions paired with SoI images) while maintaining contextual coherence across generation steps. Besides, this framework natively support for subject editing via multi-round chat, effectively addressing the overfitting limitations commonly observed in diffusion-based models.
@@ 8.5 Interactive Personalization Workflow
The evolution of interactive generation systems has unlocked new frontiers for PCS, particularly through the integration of multi-round interactive generation. This capability allows users to iteratively refine and accurately define the SoI, addressing the challenge of translating vague or complex requirements into precise content generation. For instance, conversational PCS systems like Gemini-2.0-flash [165] exemplify this progress, leveraging natural language dialogue to iteratively optimize both subject fidelity and prompt alignment. By enabling users to provide real-time feedback and adjust parameters through chat-like interactions, these systems bridge the gap between abstract intent and concrete outputs, aligning with PCS’s core objective of balancing faithful subject representation with flexible editability.
@@ 9 Conclusion
Textual Inversion [7] Object TTF LDM Token Embedding Fine-tuning
@@
DreamBooth [4] Object TTF Imagen Diffusion Model Fine-tuning; Regularization Dataset
DreamArtist [92] Object TTF LDM Negative Prompt Fine-tuning
DVAR [90] Object TTF SD 1.5 Randomness Erasing
HiPer [166] Object TTF SD Token Embedding Enhancement
P+ [18] Object TTF SD 1.4 Token Embedding Enhancement
Unet-finetune [23] Object TTF SD Parameter-efficient Fine-tuning
Jia et al. [76] Object TTF Imagen Mask-assisted Generation; Regularization
COTI [83] Object TTF SD 2.0 Data Augmentation
Gradient-Free TI [94] Object TTF SD Evolutionary Strategy
PerSAM [44] Object TTF SD Mask-assisted Generation
DisenBooth [77] Object TTF SD 2.1 Regularization
NeTI [30] Object TTF SD 1.4 Token Embedding Enhancement
Prospect [31] Object TTF SD 1.4 Token Embedding Enhancement
@@
Break-a-Scene [38] Object TTF SD 2.1 Attention-based Operation; Attention-based Operation; Mask-assisted Generation
COMCAT [96] Object TTF SD 1.4 Attention-based Operation
ViCo [71] Object TTF SD Attention-based Operation; Regularization
OFT [167] Object TTF SD 1.5 Finetuning Strategy
LyCORIS [36] Object TTF SD 1.5 Attention-based Operation
He et al. [40] Object TTF SD, SDXL Data Augmentation
DIFFNAT [168] Object TTF SD Regularization
MATTE [91] Object; Style; High-level Semantic TTF SD Regularization
LEGO [116] Object TTF SD Regularization
Catversion [95] Object TTF SD 1.5 Embedding Concatenation
@@
CLiC [45] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation
@@
HiFi Tuner [32] Object TTF SD 1.4 Attention-based Operation; Mask-assisted Generation; Regularization
InstructBooth [93] Object TTF SD 1.5 Reinforcement Learning
DETEX [49] Object TTF SD 1.4 Mask-assisted Generation
DreamDistribution [169] Object TTF SD 2.1 Token Embedding Enhancement
DreamTuner [39] Object TTF SD Attention-based Operation
Xie et al. [34] Object TTF SD Regularization; Mask-assisted Generation
Infusion [170] Object TTF SD 1.5 Attention-based Operation
Pair Customization [171] Object TTF SDXL Disentanglement Approach
DisenDreamer [172] Object TTF SD 2.1 Disentanglement Approach
DreamBooth+++++ + [173] Object TTF SD Attention-based Operation
CoRe [174] Object TTF SD 2.1 Regularization
DCI [175] Object TTF SD Token Embedding Optimization
Re-Imagen [97] Object PTA Imagen Retrieval-augmented Paradigm
Versatile Diffusion [176] Object PTA SD 1.4 Architecture Design
Tuning-Encoder [88] Object PTA SD Regularization
@@
ELITE [47] Object PTA SD 1.4 Attention-based Operation; Mask-assisted Generation
UMM-Diffusion [61] Object PTA SD 1.5 Reference Feature Injection
SuTI [52] Object PTA Imagen Data Augmentation
InstantBooth [98] Object PTA SD 1.4 Patch Feature Extraction
@@
BLIP-Diffusion [57] Object PTA SD 1.5 Data Augmentation; Mask-assisted Generation
Domain-Agnostic [89] Object PTA SD Regularization; Attention-based Operation
IP-Adapter [51] Object PTA SD 1.5 Reference Feature Injection
Kosmos-G [60] Object; Face PTA SD 1.5 Multimodal Language Modeling
Kim et al. [177] Object PTA SD 1.4 Reference Feature Injection
CAFÉ [58] Object; Face PTA SD 2.0 Multimodal Language Modeling
SAG [178] Object; Face; Style PTA SD 1.5 Gradient-based Guidance
BootPIG [179] Object PTA SD Data Augmentation
CustomContrast [180] Object PTA SD 1.5 Contrastive Learning
MoMA [59] Object PTA SD 1.5 Multimodal LLM Adapter
@@
Diptych Prompting [101] Object; Style PTA FLUX Attention-based Operation; Mask-assisted Generation
StyleDrop [24] Style; Multiple Subjects TTF Muse Data Augmentation
StyleBoost [84] Style TTF SD 1.5 Regularization Dataset
StyleAligned [99] Style TTF SDXL Regularization
GAL [19] Style; Object TTF SD 1.5 Data Augmentation
StyleForge [181] Style TTF SD 1.5 Parameter-efficient Fine-tuning
FineStyle [182] Style TTF Muse Mask-assisted Generation
Style-Friendly [100] Style TTF FLUX-dev; SD 3.5 SNR Sampler
StyleAdapter [82] Style PTA SD Data Augmentation; Reference Feature Fxtraction
ArtAdapter [183] Style PTA SD 1.5 Data Augmentation
ProFusion [106] Face TTF SD 2.0 Fusion Sampling
Celeb Basis [65] Face TTF SD Face Feature Representation
Banerjee et al. [184] Face TTF SD 1.4 Regularization
Magicapture [37] Face; Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation; Identity Loss
Concept-centric [163] Face TTF SD 1.5 Modified Classifier-free Guidance
Cross Initialization [33] Face TTF SD 2.1 Regularization
@@
OMG [27] Face; Multiple Subjects TTF SDXL Attention-based Operation; Mask-assisted Generation; Identity Loss
@@
InstantFamily [185] Face; Multiple Subjects; Extra Conditions TTF SD 1.5 Mask-assisted Generation
PersonaMagic [186] Face TTF SD 1.4 Attention-based Operation
HyperDreamBooth [26] Face TTF; PTA SD 1.5 Pretraining and Fast Fine-tuning
Su et al. [187] Face PTA LDM; StyleGAN Identity Loss; Multi-task Learning
@@
FastComposer [5] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
Face0 [107] Face PTA SD 1.4 Fusion Sampling
DreamIdentity [48] Face PTA SD 2.1 Attention-based Operation
@@
Face-Diffuser [79] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
W-Plus-Adapter [53] Face PTA SD 1.5; StyleGAN Face Feature Representation
Portrait Diffusion [188] Face PTA SD 1.5 Mask-assisted Generation
RetriNet [74] Face PTA SD Mask-assisted Generation
FaceStudio [109] Face PTA SD Attention-based Operation
PVA [189] Face PTA LDM Mask-assisted Generation
DemoCaricature [121] Face; Extra Conditions PTA SD 1.5 Regularization
@@
PhotoMaker [46] Face PTA SDXL Data Collection; Reference Feature Extraction
Stellar [35] Face PTA SDXL Evaluation Prompts and Metrics
@@
PortraitBooth [50] Face PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
@@
InstantID [6] Face; Extra Conditions PTA SDXL Multi-feature Injection
ID-Aligner [190] Face PTA SD 1.5; SDXL Feedback Learning
@@
MoA [191] Face; Multiple Subjects PTA SD 1.5 Attention-based Operation; Mask-assisted Generation
IDAdapter [192] Face PTA SD 2.1 Mixed Facial Features
Infinite-ID [193] Face PTA SDXL Reference Feature Injection
Face2Diffusion [162] Face PTA SD Multi-feature Injection
FreeCure [194] Face PTA SD 1.5; SDXL Mask-assisted Generation
Omni-ID [195] Face PTA FLUX Face representation enhancement
Custom Diffusion [20] Multiple Subjects TTF SD 1.4 Parameter-efficient Fine-tuning; Constrained Optimization
Cones [111] Multiple Subjects TTF SD 1.4 Concept Neurons Activation
SVDiff [22] Multiple Subjects TTF SD Data Augmentation; Attention-based Operation
Perfusion [21] Multiple Subjects TTF SD 1.5 Regularization
Mix-of-Show [25] Multiple Subjects TTF SD 1.5 Attention-based Operation; Mask-assisted Generation
Cones-2 [70] Multiple Subjects TTF SD 2.1 Attention-based Operation; Regularization; Mask-assisted Generation
PACGen [75] Multiple Subjects TTF SD 1.4 Mask-assisted Generation
Compositional Inversion [87] Multiple Subjects TTF SD Attention-based Operation; Regularization
EM-style [72] Multiple Subjects TTF SD Attention-based Operation; Mask-assisted Generation
MC2 [196] Multiple Subjects TTF SD 1.5 Attention-based Operation
MultiBooth [64] Multiple Subjects TTF SD 1.5 Mask-assisted Generation
Matsuda et al. [197] Multiple Subjects TTF LDM Mask-assisted Generation
MagicTailor [198] Multiple Subjects TTF SD 2.1 Mask-assisted Generation; Attention-based Operation
AnyDoor [78] Multiple Subjects PTA SD 2.1 Mask-assisted Generation
Subject-Diffusion [63] Multiple Subjects PTA SD 2.0 Attention-based Operation; Data Augmentation; Mask-assisted Generation
CustomNet [113] Multiple Subjects; Extra Conditions PTA SD Multi-condition Integration
MIGC [114] Multiple Subjects PTA SD Attention-based Operation; Mask-assisted Generation
Emu2 [199] Multiple Subjects PTA SDXL In-Context Learning
SSR-Encoder [55] Multiple Subjects PTA SD 1.5 Attention-based Operation
$\lambda$-eclipse [112] Multiple Subjects PTA Kandinsky v2.2 Contrasstive Learning
MS-Diffusion [56] Multiple Subjects PTA SD Attention-based Operation
ReVersion [115] High-level Semantic TTF SD Regularization
Inv-ReVersion [200] High-level Semantic TTF SD 1.5 Regularization
CusConcept [201] High-level Semantic TTF SD 2.1 Regularization
ADI [117] High-level Semantic TTF SD 2.1 Mask-assisted Generation
PhotoSwap [118] Extra Conditions TTF SD 2.1 Attention-based Operation
PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
Layout-Control [69] Extra Conditions TTF SD Attention-based Operation
SwapAnything [203] Extra Conditions TTF SD 2.1 Mask-assisted Generation
PE-VITON [202] Extra Conditions TTF SD Shape and Texture Control
Viewpoint Control [122] Extra Conditions TTF SDXL 3D Feature Incorporation
Prompt-Free Diffusion [204] Extra Conditions PTA SD 2.0 Reference Feature Injection
Uni-ControlNet [205] Extra Conditions PTA SD Multi-feature Injection
ViscoNet [206] Extra Conditions PTA SD Multi-feature Injection
Context Diffusion [207] Extra Conditions PTA LDM Multi-feature Injection
FreeControl [208] Extra Conditions PTA SD 1.5; SD 2.1; SDXL Fusion Guidance
Li et al. [209] Extra Conditions PTA SD Attention-based Operation; Mask-assisted Generation
@@
Tune-A-Video [127] Video TTF SD 1.4 Parameter-efficient Fine-tuning
Gen-1 [130] Video TTF LDM Multi-feature Injection
Make-A-Protagonist [210] Video TTF LDM Attention-based Operation
Animate-A-Story [131] Video TTF SD Retrieval-augmented Paradigm
MotionDirector [134] Video TTF ZeroScope; ModelScope Disentanglement Approach
LAMP [132] Video TTF SDXL; SD 1.4 First-frame Conditioned Pipeline
VMC [129] Video TTF Show-1 Parameter-efficient Fine-tuning
SAVE [128] Video TTF VDM Attention-based Operation
Customizing Motion [133] Video TTF ZeroScope Regularization
DreamVideo [135] Video TTF ModelScope Disentanglement Approach
MotionCrafter [211] Video TTF VDM Disentanglement Approach
Customize-A-Video [212] Video TTF ModelScope Multi-stage Refinement
Magic-Me [213] Video TTF SD 1.5; AnimateDiff Multi-stage Refinement
CustomTTT [214] Video TTF VDM Fine-tuning and Alignment
PersonalVideo [215] Video; Face TTF AnimateDiff Identity Injection
@@
CustomVideo [81] Video; Multiple Subjects TTF ZeroScope Data Augmentation; Attention-based Operation
VideoDreamer [80] Video; Multiple Subjects TTF SD 2.1 Data Augmentation
VideoAssembler [125] Video PTA VidRD Reference Feature Injection
VideoBooth [124] Video PTA VDM Reference Feature Injection
VideoMaker [216] Video PTA VDM Reference Feature Injection
SUGAR [217] Video PTA VDM Attention-based Operation
StyleCrafter [54] Video; Style PTA SD 1.5 Reference Feature Injection
ID-Animator [218] Video; Face PTA AnimateDiff Reference Feature Injection
ConsisID [219] Video; Face PTA CogVideoX frequency decomposition
@@
MotionCharacter [220] Video; Face PTA VDM ID-Preserving Optimization; Motion Control Enhancement
@@
DreaMoving [126] Video; Face; Extra Conditions PTA SD Multi-feature Injection
Magic3D [138] 3D TTF Imagen; LDM Score Distillation Sampling
DreamBooth3D [140] 3D TTF Imagen Multi-stage Refinement
PAS [145] 3D PTA SD Text-to-3D-Pose
StyleAvatar3D [146] 3D PTA LDM Multi-view Alignment
AvatarBooth [147] 3D TTF SD Dual Model Fine-tuning
MVDream [139] 3D TTF SD 2.1 Score Distillation Sampling
Consist3D [141] 3D TTF SD Token Embedding Enhancement
Animate124 [143] 3D TTF SD 1.5 Multi-stage Refinement
Dream-in-4D [144] 3D TTF SD 2.1 Multi-stage Refinement
TextureDreamer [142] 3D TTF LDM Texture Extraction
TIP-Editor [221] 3D TTF SD Multi-stage Refinement; Mask-assisted Generation
Anti-DreamBooth [148] Attack and Defense TTF SD 2.1 Perturbation Learning
Concept Censorship [149] Attack and Defense TTF SD 1.4 Trigger Injection
Huang et al. [222] Attack and Defense TTF SD backdoor attack
Continual Diffusion [223] Others TTF SD Continual Learning
SVGCustomization [150] Others TTF SD 1.5 Fine-tuning and Alignment
StitchDiffusion [151] Others TTF LDM Parameter-efficient Fine-tuning
MC-TI [224] Others TTF SD 1.5 Regularization